# Hepato-LLaVA: An Expert MLLM with Sparse Topo-Pack Attention for Hepatocellular Pathology Analysis on Whole Slide Images

### v1

@@ 1 Introduction
A key design choice in these MLLMs is how to represent gigapixel WSIs for MLLMs. Existing methods typically fall into two categories: thumbnail-based approaches [9] [16] that resize the gigapixel image into a megapixel image, and slide-encoder-based approaches [2] [7] that aggregate thousands of patches into global tokens. However, thumbnail-based approaches inevitably lose patch-level details, which are critical as pathologists rely on high-resolution views for diagnosis. Focusing on slide-encoder-based approaches, we investigate two primary research questions regarding the slide encoder, which acts as a bottleneck:
@@
To address RQ1, we design a Sparse Topo-Pack Attention mechanism that simulates the local aggregation and global juxtaposition of pathological diagnosis. By modeling 2D tissue topology, the slide encoder extracts topology-informed representations while significantly reducing spatial redundancy. A light-weight connector then condenses these features into fixed-length query tokens, effectively bridging the modality gap without losing critical diagnostic details. To address RQ2, we leverage hierarchical features to develop HepatoPathoVQA, a comprehensive multi-scale VQA dataset specifically for HCC. This dataset contains over 33K expert-validated QA pairs covering the full clinical workflow across three distinct scales: WSIs, Region of Interest (ROI, 2 mag.), and Patch (10 and 20 mag.). Finally, through Low-Rank Adaptation (LoRA) fine-tuning on this unified framework, we present Hepato-LLaVA, a specialized MLLM optimized for fine-grained hepatocellular pathology analysis.
@@ 2.4 Hepato-LLaVA
Multi-scale Instruction Tuning. To effectively bridge the modality gap between the pre-trained, frozen slide encoder and the MLLM, we introduce a Q-Former [6] Connector. To handle variable-length pack summary token sequences, this lightweight module employs a set of learnable query vectors to interact with inputs. The instruction tuning process is further divided into two sub-phases: (1) Visual-Language Alignment: We first train the connector on the HepatoPath-Caption dataset while keeping both the slide encoder and LLM frozen. This stage aligns visual representations with linguistic semantics, enabling the connector to translate pathological features. (2) Diagnostic Instruction Tuning: We fine-tune both the connector and the LLM on the HepatoPath-VQA dataset. This stage optimizes the model for diagnostic inference, enabling it to interpret multi-scale visual evidence and generate clinically precise responses.
@@ 3.2 Main Results
Hepato-LLaVA is initialized from WSI-LLaVA, which is fine-tuned using LoRA (r=128), while the connector is trained from scratch. We evaluated Hepato-LLaVA on HepatoPathoBench against six baselines. As shown in Table 1, while Thumbnail-based models yielded suboptimal results (Avg 0.50–0.57) due to severe information loss, Hepato-LLaVA achieved state-of-the-art performance (Avg 0.83). This establishes a substantial lead of 0.17 over the best WSI-based runner-up SlideChat (0.66). In open-ended tasks, Hepato-LLaVA achieved a WSI-P of 0.79 in morphology and 0.75 in diagnosis, surpassing the SlideChat (0.70, 0.72). In close-ended tasks, our model attained 0.97 accuracy in morphological single-choice and 0.88 in multi-choice, significantly outperforming Patho-R1 (0.87, 0.50), a strong baseline enhanced by reinforcement learning. Regarding multi-scale consistency, Hepato-LLaVA demonstrated robust performance across WSI (0.82), ROI (0.83), and Patch (0.83) levels. This represents an improvement over our backbone model WSI-LLaVA (0.65, 0.67, 0.64), validating that the proposed sparse attention mechanism effectively overcomes the scale variance limitations inherent in the base architecture. Qualitative evidence in Fig. 3 confirms its ability to interpret critical morphological evidence for precise diagnosis.

### v2

@@ 1 Introduction
A key design choice in these MLLMs is how to represent gigapixel WSIs for MLLMs. Existing methods typically fall into two categories: thumbnail-based approaches [9] [16] that resize the gigapixel image into a megapixel image, and slide-encoder-based approaches [2] [7] that aggregate thousands of patches into global tokens. However, thumbnail-based approaches inevitably lose patch-level details, while both methods ingest only WSIs inputs, consequently less equipped with patch-level capabilities essential for diagnosis. We investigate two primary research questions regarding the slide encoder, which acts as a bottleneck:
@@
To address RQ1, we design a Sparse Topo-Pack Attention mechanism that simulates the local aggregation and global juxtaposition of pathological diagnosis. By modeling 2D tissue topology, the slide encoder extracts topology-informed representations while significantly reducing spatial redundancy. A light-weight connector then condenses these features into fixed-length query tokens, effectively bridging the modality gap without losing critical diagnostic details. To address RQ2, we leverage hierarchical features to develop HepatoPathoVQA, a comprehensive multi-scale VQA dataset specifically for HCC. This dataset contains over 33K expert-validated QA pairs covering the full clinical workflow across three distinct scales: WSIs, Region of Interest (ROI, 5 mag.), and Patch (10 and 20 mag.). Finally, through Low-Rank Adaptation (LoRA) fine-tuning on this unified framework, we present Hepato-LLaVA, a specialized MLLM optimized for fine-grained hepatocellular pathology analysis.
@@ 2.4 Hepato-LLaVA
Multi-scale Instruction Tuning. To effectively bridge the modality gap between the pre-trained, frozen slide encoder and the MLLM, we introduce a Q-Former [6] Connector. To handle variable-length pack summary token sequences, this lightweight module employs a set of learnable query vectors to interact with inputs. The instruction tuning process is further divided into two sub-phases: (1) Visual-Language Alignment: We first train the connector on the HepatoPath-Caption dataset while keeping both the slide encoder and MLLM frozen. This stage aligns visual representations with linguistic semantics, enabling the connector to translate pathological features. (2) Diagnostic Instruction Tuning: We fine-tune both the connector and the MLLM on the HepatoPath-VQA dataset. This stage optimizes the model for diagnostic inference, enabling it to interpret multi-scale visual evidence and generate clinically precise responses.
@@ 3.2 Main Results
Hepato-LLaVA is initialized from WSI-LLaVA, which is fine-tuned using LoRA (r=128), while the connector is trained from scratch. We evaluated Hepato-LLaVA on HepatoPathoBench against six baselines. As shown in Table 1, while Thumbnail-based models yielded suboptimal results (Avg 0.50–0.57) due to severe information loss, Hepato-LLaVA achieved state-of-the-art performance (Avg 0.83). This establishes a substantial lead of 0.17 over the best WSI-based runner-up SlideChat (0.66). In open-ended tasks, Hepato-LLaVA achieved a WSI-P of 0.79 in morphology and 0.75 in diagnosis, surpassing the SlideChat (0.70, 0.72). In close-ended tasks, our model attained 0.97 accuracy in morphological single-choice and 0.88 in multi-choice, significantly outperforming Patho-R1 (0.87, 0.50), a strong baseline enhanced by reinforcement learning. Regarding multi-scale consistency, Hepato-LLaVA demonstrated robust performance across WSI (0.82), ROI (0.83), and Patch (0.83) levels. This represents an improvement over our backbone model WSI-LLaVA (0.65, 0.67, 0.64), validating that the proposed sparse attention and connector effectively overcomes the scale variance limitations inherent in the base architecture. Qualitative evidence in Fig. 3 confirms its ability to interpret critical morphological evidence for precise diagnosis.